The SVJour document class users guide supplement for global use

\begin{matrix}c \\ ◯ \end{matrix} 1998, Springer Verlag Heidelberg All rights reserved.

24 April 1998

1 Introduction

This document describes the global option for the SVJour L^{A}TEX 2ε document class. It is meant for journals (of one- or two-column style) that have not yet been supported via a particular class option.

For details on manuscript handling and the review process we refer to the Instructions for authors in the printed journal you are writing your article for. For style matters please consult also previous issues of “your” journal.

2 Initializing the Class

Please make sure that no particular class option exists for your current project, take a good look at the announcing web page

       http://www.springer.de/author/tex/help-journals.html

If the journal is not listed there the specific package is not yet ready and you should begin your document with

   \documentclass[global]{svjour}

This holds for a one-column journal – for a two-column journal just add the option twocolumn so that the documentclass command now reads

   \documentclass[global,twocolumn]{svjour}

All other options are also described in the main Users guide.

3 Things to Note

There may be a slightly different layout for the journal you plan to create an article for and which is not taken care for by the global option. This is not a great problem since all the articles are reprocessed anyway to include final page numbers, bibliographic information, and so on. The final layout is applied to your article during this process thus deriving benefit from the general use of L^{A}TEX 2ε and the SVJour document class altogether.

Since Springer-Verlag publishes over 400 different journals – over 50 thereof with TEX – it may take a while to produce a specific TEX macro package just for the one you are writing for at the moment. Please keep an eye on the web page mentioned above.